Solve the following sytem by substitution:
4x - 12y = 5
-x + 3y = -1
Thanks

Answer by tutorcecilia(2152)

4x - 12y = 5
-x + 3y = -1
.
Solve for one of the variables and substitute the solution into the other equation. solve for the second variable.
-x+3y=-1 [Solve for x (or solve for y)]
-x=-3y-1
x=3y+1 [substitute this solution into the other equation and solve for y]
.
4x-12y=5
4(3y+1)-12y=5
12y+4-12y=5
4 does not equal to 5, therefore there is no solution to this system. This means that the lines never intersect or cross.